# High-order anisotropic flow and nonlinear hydrodynamic responses

### v1

@@ I Introduction
$\mathcal{E}_{n}=\varepsilon_{n}e^{in\Psi_{n}^{\mathrm{PP}}}=-\frac{\int d^{2}re(r,\phi)e^{in\phi}}{\int d^{2}re(r,\phi)},$ (2)
where the weight $e(r,\phi)$ is the initial energy density profile. For harmonic order $n=2,3$, the anisotropic flow coefficients show strong. linear responses to the collision geometry, $V_{n}=k_{n}\mathcal{E}_{n}$. The high-order anisotropic flow coefficient $V_{n}$ ($n>3$) not only exhibits dependence on the $n$-th order eccentricity $\mathcal{E}_{n}$ of the initial density profile, probing fine details about the medium at small length scales, it also receives contributions from the low-order deformation via nonlinear interactions during the hydrodynamic evolution, serving as a potential probe for the QGP viscosity and speed of sound in the equation of state for dense nuclear matter [21, 22].
The high statistics runs from the LHC experiments have recently pushed flow correlation measurements to multi-particles, such as $v_{2}\{10\}$ [23], and high-order flow harmonics. The ALICE Collaboration has recently measured two-particle charged hadron anisotropic flow coefficients $v_{n}\{2\}$ up to $n=9$ in Pb-Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV [24]. In this paper, we will perform high-statistics numerical simulations to make theoretical comparisons with these measurements. We will quantify the relative contributions of linear and nonlinear modes to the high-order $v_{n}$ for $n=4-9$, providing insight into these measurements.
@@ II Linear and nonlinear anisotropic flow responses
We can generally decompose the complex high-order anisotropic flow coefficients $V_{n}$ ($n\geq 4$) into linear and nonlinear terms, where the linear term represents a response to the initial geometry deformation at the same order ($\mathcal{E}_{n}$). The nonlinear modes arise from fluctuations in lower-order harmonics ($m<n$[25]. Let us start with the decomposition of $V_{4}$ and $V_{5}$ as follows,
@@
Similarly, based on Eqs. (11), (12), and (13), we can construct a $3\times 3$ matrix equation for $V_{7}$, a $6\times 6$ matrix equation for $V_{8}$, and a $7\times 7$ matrix equation for $V_{9}$, respectively.
@@ III Phenomenological study at the LHC
In every hydrodynamic event, we compute two sub-event anisotropic flow vectors, namely $\{V_{n}^{A}\}$ with kinematic cuts $p_{T}\in[0.2,3]$ GeV and $\eta\in[-3.2,-0.4]$ and $\{V_{n}^{B}\}$ using charged hadrons with $p_{T}\in[0.2,3]$ GeV and $\eta\in[0.4,3.2]$. We compute $v_{n}\{2\}$ as
$v_{n}\{2\}$ =$=\sqrt{\Re\{\langle V_{n}^{A}(V_{n}^{B})^{*}\rangle\}}$
@@ III.1 The anisotropic flow power spectrum
The power spectrum of temperature fluctuations is a powerful tool for studying cosmic microwave background (CMB) radiation [36]. In central Pb+Pb collisions, all the anisotropic harmonic flow coefficients are driven by event-by-event fluctuations. Ref. [37] proposed that $v_{n}\{2\}$ as a power spectrum has the dependence $v_{n}\{2\}\propto\exp[-bn^{2}]$, where the coefficient $b$ depends on the specific viscosity used in the hydrodynamic simulations.
@@
Figure 1 shows the $v_{n}\{2\}$ power spectrum in 0-5% Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. In addition to the default simulations with temperature-dependent shear and bulk viscosity calibrated in Ref. [32], we also show a reference result using a constant $\eta/s=0.06$ and $\zeta/s=0$ for comparison. The power spectrum in Fig. 1 shows that the high-order $v_{n}\{2\}$ coefficients have strong sensitivity to the viscosity used in the simulations. For both cases, the anisotropic flow coefficients $v_{2}\{2\}$ to $v_{6}\{2\}$ show the $\exp[-bn^{2}]$ dependence and the coefficient $b$ is smaller for the $\eta/s=0.06,\zeta/s=0$ case. The high-order $v_{n}\{2\}$ with $n\geq 7$ decays slower than $e^{-n^{2}}$.
Figure 2: Charged hadron anisotropic flow coefficients $v_{n}\{2\}(n=2-9)$ in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV compared with the ALICE measurements [24]. Individual panels present the results in different centrality bins from central up to 50%.
Figure 3: The charged hadron $v_{n}\{2\}$ ratios for data over theory as functions of the harmonic order from Fig. 2.
Figure 4: Model prediction for charged hadron $v_{n}\{2\}$ as a function of $n$ $(n=1-12)$ in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.36$ TeV.
Figure 2 shows the comparison with the ALICE measurement on anisotropic flow coefficients $v_{n}\{2\}$ as functions of the harmonic order $n$ for six centrality bins in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. Figure 3 shows the ratios between the experimental measurements and the theoretical calculations. We note that the model calibration performed in Ref. [32] only focused on $v_{2}\{2\}$ measurements. Remarkably, the IP-Glasma initial conditions can provide a simultaneous and accurate description of the ALICE measurements from $v_{2}\{2\}$ up to $v_{7}\{2\}$ over all the centrality bins. This comparison demonstrates that the IP-Glasma model captures initial-state fluctuations over a wide range of scales [38]. Looking at higher-order flow coefficients, we observe that the ALICE measured $v_{8}\{2\}$ and $v_{9}\{2\}$ saturate around $1\times 10^{-3}$, while our theoretical results continue to decay exponentially.
@@ IV Conclusion
We also performed systematic nonlinear mode decomposition for high-order anisotropic flow coefficients up to the harmonic order $n=9$. Based on our model calculations, solving the full matrix equation is crucial for obtaining the correct values of some nonlinear mode coefficients for high-order anisotropic flow coefficients with harmonic order $n\geq 6$. We provide model predictions for the nonlinear mode coefficients for $V_{8}$ and $V_{9}$. We further compare the RMS of the linear flow components $\sqrt{\langle|V_{nL}|^{2}\rangle}$ with $v_{n}\{2\}$. In semi-peripheral collisions, we find an interesting non-monotonic dependence of the ratios to $v_{n}\{2\}$ as a function of the harmonic order $n$ with the calibrated QGP viscosity in the simulations. The maximum mixture between the linear and nonlinear modes happens at $n=6$. The nonlinear modes for $v_{n}$ with $n\geq 7$ are strongly damped by the viscosity compared to their linear modes. This finding is verified through simulations with low QGP viscosity, which show a significant nonlinear mode mixture for all high-order anisotropic flow coefficients. It will be interesting to search for these observations in the upcoming high-statistics experimental measurements.
@@ Appendix A Numerical convergence of $v_{n}\{2\}$
Because hydrodynamic simulations utilize rectangular-shaped grids in the transverse plane, we perform convergence tests by rotating every collision event with random angles, ensuring that the rectangular grid is not aligned with any specific axis. Figure 14 shows that the relative variation is below 1% for $v_{n}\{2\}$ with $n\leq 6$. The relative difference increases to $\sim 5\%$ for $v_{9}\{2\}$.
@@ Appendix B Decoupling of the directed flow $V_{1}$
Based on the results in Fig. 15, we can replace $V_{2L}$ and $V_{3L}$ by $V_{2}$ and $V_{3}$, respectively. Similar to Eq. (15), we can solve the $2\times 2$ matrix equations for the nonlinear mode coefficients for $V_{4}$ and $V_{5}$ when the $V_{1}$ terms are present.

### v2

@@ I Introduction
$\mathcal{E}_{n}=-\frac{\int d^{2}re(r,\phi)e^{in\phi}}{\int d^{2}re(r,\phi)},$ (2)
where the weight $e(r,\phi)$ is the initial energy density profile. For harmonic order $n=2,3$, the anisotropic flow coefficients show strong linear responses to the collision geometry, $V_{n}=k_{n}\mathcal{E}_{n}$. The high-order anisotropic flow coefficient $V_{n}$ ($n>3$) not only exhibits dependence on the $n$-th order eccentricity $\mathcal{E}_{n}$ of the initial density profile, probing fine details about the medium at small length scales, it also receives contributions from the low-order deformation via nonlinear interactions during the hydrodynamic evolution, serving as a potential probe for the QGP viscosity and speed of sound in the equation of state for dense nuclear matter [21, 22].
The high statistics runs from the LHC experiments have recently pushed flow correlation measurements to multi-particles, such as $v_{2}\{10\}$ [23, 24], and high-order flow harmonics. The ALICE Collaboration has recently measured two-particle charged hadron anisotropic flow coefficients $v_{n}\{2\}$ up to $n=9$ in Pb-Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV [25]. In this paper, we will perform high-statistics numerical simulations to make theoretical comparisons with these measurements. We will quantify the relative contributions of linear and nonlinear modes to the high-order $v_{n}$ for n=49n=4-9, providing insight into these measurements.
@@ II Linear and nonlinear anisotropic flow responses
We can generally decompose the complex high-order anisotropic flow coefficients $V_{n}$ ($n\geq 4$) into linear and nonlinear terms, where the linear term represents a response to the initial geometry deformation at the same order ($\mathcal{E}_{n}$). The nonlinear modes arise from fluctuations in lower-order harmonics (m<nm<n[26]. Let us start with the decomposition of $V_{4}$ and $V_{5}$ as follows,
@@
Similarly, based on Eqs. (11), (12), and (13), we can construct matrix equations for $V_{7}$ to $V_{9}$.
@@ III Phenomenological study at the LHC
In every hydrodynamic event, we compute two sub-event anisotropic flow vectors, namely $\{V_{n}^{A}\}$ with kinematic cuts $p_{T}\in[0.2,3]$ GeV and $\eta\in[-3.2,-0.4]$ and $\{V_{n}^{B}\}$ using charged hadrons with $p_{T}\in[0.2,3]$ GeV and $\eta\in[0.4,3.2]$. We note that the ALICE experiment measured $v_{n}\{2\}$ with $p_{T}\in[0.2,5]$ GeV. We have verified that the relative difference of $v_{n}\{2\}$ is within 4% of our model’s prediction between the two $p_{T}$ cuts, within our statistical uncertainties. We compute $v_{n}\{2\}$ as
$v_{n}\{2\}$ $=\sqrt{\Re\{\langle V_{n}^{A}(V_{n}^{B})^{*}\rangle\}}$
@@ III.1 The anisotropic flow power spectrum
The power spectrum of temperature fluctuations is a powerful tool for studying cosmic microwave background (CMB) radiation [37]. In central heavy-ion collisions, all the anisotropic harmonic flow coefficients are driven by event-by-event fluctuations, where all orders of initial eccentricity $|\mathcal{E}_{n}|$ are comparable in their sizes [7]. Refs. [38, 21, 39, 40] proposed that the eccentricity scaled anisotropic flow $v_{n}/|\mathcal{E}_{n}|\propto\exp[-bn^{2}]$ with the coefficient $b$ depending on the specific viscosity used in the hydrodynamic simulations. In contrast, Ref. [41, 42] showed that the $v_{n}/|\mathcal{E}_{n}|$ decays exponentially at large $n$.
@@
Figure 1 shows the $v_{n}\{2\}$ power spectrum in 0-5% Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. In addition to the default simulations with temperature-dependent shear and bulk viscosity calibrated in Ref. [33], we also show a reference result using a constant $\eta/s=0.06$ and $\zeta/s=0$ for comparison. The power spectrum in Fig. 1 shows that the high-order $v_{n}\{2\}$ coefficients have strong sensitivity to the viscosity used in the simulations.
In central Pb+Pb collisions, we can assume the magnitudes of initial eccentricity $|\mathcal{E}_{n}|$ are roughly the same for all order $n$ [7]. The anisotropic flow coefficients $v_{2}\{2\}$ to $v_{6}\{2\}$ from our model show the $\exp[-bn^{2}]$ dependence for both cases, with the coefficient $b$ is smaller for the $\eta/s=0.06,\zeta/s=0$ case, qualitatively agreeing with the proposed relation from Refs. [38, 21, 39, 40]. In the meantime, the high-order $v_{n}\{2\}$ with $n\geq 7$ decays exponentially with $n$, consistent with the expectations from Refs. [41, 42].
In the following studies, we present results using temperature-dependent shear and bulk viscosity, calibrated in Ref. [33].
Figure 2: Charged hadron anisotropic flow coefficients $v_{n}\{2\}(n=2-9)$ in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV with two kinematic cuts from the boost-invariant simulations compared with the ALICE measurements [25]. Individual panels present the results in different centrality bins from central up to 50%.
Figure 3: The charged hadron $v_{n}\{2\}$ ratios for the ALICE data over the theory simulations as functions of the harmonic order from Fig. 2.
Figure 4: Model prediction for charged hadron $v_{n}\{2\}$ as a function of $n$ (n=112)(n=1-12) in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.36$ TeV.
Figure 2 shows the comparison with the ALICE measurement on anisotropic flow coefficients $v_{n}\{2\}$ as functions of the harmonic order $n$ for six centrality bins in Pb+Pb collisions at $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV. Our boost-invariant simulations yield compatible $v_{n}\{2\}$ results with the different kinematic cuts. The results with the large rapidity acceptance have smaller statistical uncertainties. Figure 3 shows the ratios between the experimental measurements and the theoretical calculations. We note that the model calibration performed in Ref. [33] only focused on $v_{2}\{2\}$ measurements. Remarkably, the IP-Glasma initial conditions can provide a simultaneous and accurate description of the ALICE measurements from $v_{2}\{2\}$ to $v_{7}\{2\}$ across all centrality bins. This comparison demonstrates that the IP-Glasma model captures initial-state fluctuations over a wide range of scales [43]. Looking at higher-order flow coefficients, we observe that the ALICE measured $v_{8}\{2\}$ and $v_{9}\{2\}$ saturate around $10^{-3}$, while our theoretical results continue to decay exponentially.
@@ III.2 Anisotropic flow mode decomposition for $v_{4}-v_{9}$
Compared to the previous IP-Glasma + MUSIC results in Refs. [36, 25], our current model provides a better overall description of the non-linear mode coefficients because the current initial-state model includes sub-nucleonic structures constrained by the HERA diffractive $J/\Psi$ measurements, and it leads to different optimal QGP-specific shear and bulk viscosities [33].
@@ IV Conclusion
We also performed systematic nonlinear mode decomposition for high-order anisotropic flow coefficients up to the harmonic order $n=9$. Based on our model calculations, solving the full matrix equation is crucial for obtaining the correct values of some nonlinear mode coefficients for high-order anisotropic flow coefficients with harmonic order $n\geq 6$. It would be great to verify this difference in future experimental measurements. We provide model predictions for the nonlinear mode coefficients for $V_{8}$ and $V_{9}$. We further compare the RMS of the linear flow components $\sqrt{\langle|V_{nL}|^{2}\rangle}$ with $v_{n}\{2\}$. In semi-peripheral collisions, we find an interesting non-monotonic dependence of the ratios to $v_{n}\{2\}$ as a function of the harmonic order $n$ with the calibrated QGP viscosity in the simulations. The maximum mixture between the linear and nonlinear modes happens at $n=6$. The nonlinear modes for $v_{n}$ with $n\geq 7$ are strongly damped by the viscosity compared to their linear modes. This finding is verified through simulations with low QGP viscosity, which show a significant nonlinear mode mixture for all high-order anisotropic flow coefficients. It will be interesting to search for these observations in the upcoming high-statistics experimental measurements.
The data that support the findings of this article are openly available [44], embargo periods may apply.
@@ Appendix A Numerical convergence of $v_{n}\{2\}$
Because hydrodynamic simulations utilize rectangular-shaped grids in the transverse plane, we perform convergence tests by rotating every collision event with random angles, ensuring that the rectangular grid is not aligned with any specific axis. Figure 14 shows that the relative variation is below 1% for $v_{n}\{2\}$ with $n\leq 6$. The relative difference increases to $\approx 5\%$ for $v_{9}\{2\}$.
@@ Appendix B Decoupling of the directed flow $V_{1}$
Based on the results in Fig. 15, we can replace $V_{2L}$ and $V_{3L}$ by $V_{2}$ and $V_{3}$, respectively. Similar to Eq. (15), we can solve the matrix equations for the nonlinear mode coefficients for $V_{4}$ and $V_{5}$ when the $V_{1}$ terms are present.